# The Impact of Surface Co-location and Eye-tracking on Mixed Reality Typing

### v2

@@ 4.2. Procedure
Participants used a HoloLens 2 headset while seated. The keyboard was placed midair in front of the participant. Participants were instructed to adjust the height of the chair such that they could comfortably reach the keys. Since the keyboard was intended to be used in midair only, no table was included in Experiment 2.
The study took approximately one hour. Participants first completed an initial questionnaire and then proceeded to their first condition. Between conditions, participants took a two-minute break and completed a questionnaire about the condition just finished. During this time, we instructed them to remove the headset and walk around the room to reduce eyestrain. After all conditions, participants completed a final questionnaire. The questionnaires were identical to Experiment 1 and are included in our supplementary materials.333https://osf.io/q4n72 Participants received a $15 USD gift card.
The order of conditions was completely counterbalanced between participants. Each condition consisted of two practice sentences, followed by eight evaluation sentences. We only analyzed the evaluation sentences. For each condition, sentences were pulled randomly without replacement from the “mem1-5” set from the mobile Enron dataset (Vertanen and Kristensson, 2011). These sentences were in mixed case and included punctuation and sometimes numbers. Participants were instructed to type using both hands.
24 participants between the ages of 18 and 22 completed Experiment 2. Of these, 17 were male, 5 were female, and 2 preferred not to say. 23 participants strongly agreed with “I consider myself a fluent speaker of English”, and the remaining participant agreed with the statement. A total of seven participants reported AR experience, with four participants having used a HoloLens before, and three participants reporting use of phone AR apps (e.g. Pokémon Go).
@@ 4.3. Results
We measured entry rate as in Experiment 1. Entry rates were 10.9 in Index, 9.4 in IndexEye, 7.1 in Ten, and 7.4 WPM in TenEye (Table 3 left and Figure 8 left). This difference was significant. All post-hoc pairwise comparisons between conditions were significant except between Ten and TenEye.
Index | 10.9 1.4 [8.5, 13.9] | 0.63 0.82 [0.00, 3.11] | 0.12 0.06 [0.01, 0.27] |
IndexEye | 9.4 1.3 [6.9, 11.5] | 0.58 0.92 [0.00, 2.89] | 0.08 0.07 [0.02, 0.31] |
Ten | 7.1 1.7 [3.1, 10.5] | 1.91 2.21 [0.00, 9.63] | 0.30 0.14 [0.07, 0.62] |
TenEye | 7.4 1.9 [4.3, 11.0] | 1.27 1.54 [0.00, 6.52] | 0.16 0.11 [0.04, 0.39] |
Omnibus test | , , | , , | , , |
Pairwise | Index IndexEye, | Index IndexEye, | Index IndexEye, |
post-hoc | Index Ten, | Index Ten, | Index Ten, |
Index TenEye, | Index TenEye, | Index TenEye, | |
IndexEye Ten, | IndexEye Ten, | IndexEye Ten, | |
IndexEye TenEye, | IndexEye TenEye, | IndexEye TenEye, | |
Ten TenEye, | Ten TenEye, | TenEye Ten, |
@@
As in Experiment 1, we measured typing errors using character error rate (CER). The CER was 0.63% for Index, 0.58% for IndexEye, 1.91% for Ten, and 1.27% for TenEye (Table 3 center and Figure 8 center). While this difference was significant, the only significant post-hoc pairwise comparison was between IndexEye and Ten.
To measure error correction activity, we calculated the number of backspaces per final output character for each phrase. The number of backspaces pressed per output character was 0.12 for Index, 0.08 for IndexEye, 0.30 for Ten, and 0.16 for TenEye (Table 3 right and Figure 8 right). This difference was significant. Post-hoc pairwise comparisons were significant with the exception of between Index and IndexEye, and between Index and TenEye. Notably, eye-tracking did significantly reduce the number of corrections required in ten-finger typing. This suggests that our filtering feature was successful at preventing some erroneous keystrokes.
@@
Heatmap showing how often each participant used their left or right index finger.
Heatmap showing how often each participant used each of their fingers.
Figure 10 shows the average distance between the pressed key and eye gaze, before and after the key was pressed. In all conditions, the user looked most closely at the key at the instant it was pressed. Despite users tending to look more closely at the pressed key in the IndexEye and TenEye conditions, likely due to the restraints imposed by our eye-tracking feature, all four conditions followed the same pattern of eye movement. This supports our theory that users will naturally tend to look where they intend to type, and indicates potential for incorporating this implicit gaze signal to improve a keyboard’s interface or algorithms (e.g. auto-correct or word predictions).
We calculated how often participants used each of their fingers during all conditions. Figure 11 shows that participants in the Index and IndexEye conditions tended to share typing between their two hands. For example, 11 participants in the Index condition used their left index finger 40% of the time or higher. As shown in Figure 12, in the ten-finger conditions, in general participants used all their fingers. However, most participants showed a strong preference to typing with their index fingers. In the Ten condition, on average participants typed with their left and right index fingers 17% and 30% of the time respectively. Interestingly, in the TenEye condition, this tendency was less pronounced at 12% and 21% for the left and right index fingers respectively. This shows that the eye-tracking feature may have increased participants’ confidence in typing with other fingers.
@@ 4.4. Subjective feedback
After each condition, participants rated three statements on a 7-point Likert scale (Figure 13). Participants’ average ratings for “I entered text quickly” were: Index 4.4, IndexEye 3.9, Ten 3.1, and TenEye 3.3. There was a significant difference between ratings (), but none of the pairwise differences were significant. Participants’ average ratings for “I entered text accurately” were: Index 5.2, IndexEye 5.8, Ten 3.5, and TenEye 4.5. There was a significant difference between ratings (). Post-hoc tests showed Index was rated more accurate than Ten (), IndexEye was more accurate than Ten (), and IndexEye was more accurate than TenEye (). Other pairs were non-significant. Participants’ average ratings for “Entering text was physically easy and not tiring” were: Index 4.2, IndexEye 4.4, Ten 3.7, and TenEye 4.5. There was not a significant difference between ratings ().
@@
In a final questionnaire, participants ranked the conditions and were asked to explain their choice. 16 of 24 participants ranked Index first. Of those ranking Index the best, some representative comments were: “I didn’t have to worry about other fingers getting in the way”, “index fingers are both faster and more accurate for this”, and “I prefer the comfort of being able to relax my other fingers”. The second most preferred condition, by seven participants, was IndexEye. A representative comment from a participant preferring this condition was: “index finger tracking only was the easiest, and when combined with eye-tracking, it was also the most accurate”. Only one participant preferred TenEye.
Regarding the index finger conditions, participants seemed to prefer no eye-tracking, with 17 participants ranking Index above IndexEye. When only using index fingers, most participants seemed to think that the eye-tracking in IndexEye was more limiting than helpful. In the ten-finger conditions, however, participants seemed to prefer the eye-tracking feature, with 17 participants ranking TenEye above Ten. Participant comments specific to each condition are listed in Table 4. For all user feedback, see our supplementary materials.444https://osf.io/q4n72
@@
Condition | Comment | |
Index | + | “It was mentally easier because there was less error with less fingers able to hit random buttons.” |
+ | “More of a freedom to look around and plan for the next key to tap rather than be constrained by my eyes.” | |
+ | “The other fingers couldn’t get in the way and I didn’t have to look at the space bar to press it.” | |
- | “The press detection feels a little unreliable.” | |
- | “I did not like the constraint of index fingers only because sometimes I would use my other fingers out of habit.” | |
- | “I still was looking down most of it because I don’t usually type with only my index fingers.” | |
IndexEye | + | “The eye-tracking is really neat, also definitely made me more accurate.” |
+ | “I was able to make the fewest mistakes in this condition.” | |
+ | “Only using index fingers made it easier to type the right keys.” | |
- | “Needing to look at the keys I want to type slows me down a little.” | |
- | “The lag from the eye-tracking.” | |
- | “It took me a bit to get used to not typing with all my fingers.” | |
Ten | + | “The speed of typing with unrestricted 10 fingers was fun.” |
+ | “A bit less visually cluttered.” | |
+ | “Being able to use my thumb to press space, and pinky to press shift.” | |
- | “I got frustrated a bit so I started using fewer fingers at times because it was easier.” | |
- | “I often found myself hitting space or another key when I didn’t want to.” | |
- | “Ring finger tracking inaccuracies.” | |
TenEye | + | “Definitely more accurate, was quicker typing because of that. Most false presses weren’t registered.” |
+ | “The flexibility of using all fingers.” | |
+ | “The eye-tracking made it harder to mis-type letters on the keyboard.” | |
- | “The HoloLens had trouble reading the ring and pinky fingers from most natural angles.” | |
- | “Was difficult to type quickly at first because I was getting used to the ‘look when you type’ system.” | |
- | “Not being able to use the space bar without looking at it.” |
@@ 4.5. Discussion
Our goals in Experiment 2 were to 1) examine if users could type faster in midair using all ten fingers, and 2) to see if eye-tracking could be used to reduce accidental key presses. However, we found users had a significantly higher entry rate in the Index and IndexEye conditions than in the Ten and TenEye conditions. Users also expressed an overwhelming preference for the index-finger conditions. Users performed poorly in the Ten condition as evidenced by lower entry rates than the two index-finger conditions and more backspaces per output character then all other conditions. Tactile feedback or improved tracking might have improved performance. Participants frequently complained in open feedback about the HoloLens’ hand-tracking. They also reported difficulty pressing keys using their ring fingers and frustration with a lack of responsiveness from the HoloLens.
@@
Our hypothesis that users generally looked at the key they were attempting to press was supported by the the eye-tracking data gathered during the study. We examined this by noting how often people were looking near the key that they pressed in the Index and Ten conditions, which did not use eye-tracking to disable keys. If we had disabled keys far from their gaze point as in IndexEye and TenEye, key presses would have been excluded 32.9% and 36.0% of the time in the Index and Ten conditions respectively. This suggests that participants often looked where they were typing even when they were not required to do so. However, the IndexEye condition only excluded 10.4% of key presses. This notably lower amount may be simply because participants were more careful to match their gaze to the key they were pressing, lowering the number of unintentional key presses. However, in the TenEye condition, 36.7% of key presses occurred on disabled keys. This result is similar to the Ten condition and may show eye-tracking successfully filtered out unintentional key presses. However, it is important to note we do not know how many discarded key presses were due to unintentional presses being helpfully discarded versus key presses a participant wanted to make but that were discarded because they were not looking close enough to the key.
There is also other evidence that eye-tracking was successful. Firstly, responses indicated that eye-tracking generally did not increase eyestrain or physical fatigue experienced by users (Figure 13 bottom). Second, the addition of eye-tracking in TenEye significantly reduced the amount of backspaces per character (Figure 8 right). Lastly, character error rate was reduced in TenEye to levels similar to the index finger conditions. Overall, our results demonstrate that eye-tracking is an effective tool that should be considered when implementing ten-finger midair typing systems.
@@
A common complaint during the eye-tracking conditions was needing to look at the space bar to trigger it. Since the space bar is large and thumbs often do not move far from it when typing with ten fingers, participants felt this unnecessarily slowed typing. In the future, allowing space bar presses regardless of gaze position may improve results. Another common complaint was that the keys were unresponsive or difficult to trigger. We are uncertain if this was due to hand-tracking failures or an unnecessarily deep threshold for triggering keys. Reducing the distance that participants needed to push keys would have made it easier for participants to trigger individual keys, potentially improving performance. However, a shorter distance threshold could also increase the frequency of accidental key presses. The half-second cooldown may have also limited text entry speed in some circumstance, but given all participants in our studies typed substantially slower than 24 WPM on average (i.e. two characters per second), we think this did not unduly slow typing. Additional work could investigate if a different cooldown could improve performance. However, similar to reducing the trigger distance, this could increase typing errors and actually slow typing.
@@
Our entry rates were undoubtedly limited by the deterministic nature of our keyboard. It is likely users could have typed substantially faster if they only had to aim approximately within a key or so of their desired letters. Powerful auto-correct might even be able to sift through the many false key triggers typical of ten-finger midair typing. However, even in the case a MR keyboard supporting fast typing via auto-correct, it would still need a fallback mode to support users entering exactly what they want. This can be necessary for certain types of content, or to avoid an “auto-correct trap” in which a user’s desired and correct text is repeatedly auto “corrected” into incorrect text. Our work here can help guide how an auto-correcting MR keyboard can best support fallback to deterministic input.
An interesting avenue for future research would be to use auxiliary sensors (e.g. the AR headset’s microphone or a smartwatch) to detect taps on a physical surface instead of, or in addition to, finger collision with the keys. This could improve tap detection accuracy of virtual keyboards co-located with physical surfaces. The tactile feedback of actual contact with a surface could mitigate the problem of unintentional key presses, improving ten-finger typing.
In our midair keyboard experiments, we used a user’s gaze location to make a hard decision that deactivated keys more than 2.6 cm away. This threshold was determined based on pilot testing and given the eye-tracking accuracy of our headset. It is possible a different threshold, or a more accurate eye-tracker, would improve performance of this feature. It would also be interesting to investigate improving location or tap detection accuracy via a softer, probabilistic fusion of the hand and eye-tracking signals. We have included in our supplementary materials555https://osf.io/q4n72 detailed participant log files from both experiments. This includes when participants triggered individual keys as well as continuously sampled observations from the hand and eye tracker for Experiment 2. This data could allow exploring potential gains from incorporating such changes into future midair keyboards. We have also included the source code for the keyboard applications we used in both experiments.

### v3

@@ 4.2. Procedure
The study took approximately one hour. Participants first completed an initial questionnaire. They then donned the HoloLens 2 headset and calibrated the eye tracker before proceeding to their first condition. The keyboard was placed midair in front of the participant with the bottom of the spacebar located 40 cm below eye level. In pilot testing, we determined this to be a comfortable seated typing location for most people. Participants were instructed to adjust the chair height and location to fine-tune their position with respect to the keyboard. Since the keyboard was intended to be used in midair only, no table was used. Participants were instructed to type using both hands.
Between conditions, participants took a two-minute break and completed a questionnaire about the condition just finished. During this time, we instructed participants to remove the headset and walk around the room to reduce eyestrain. After all conditions, participants completed a final questionnaire. The questionnaires were identical to Experiment 1 and are included in our supplementary materials.555https://osf.io/q4n72 Participants received a $15 USD gift card.
The order of conditions was completely counterbalanced between participants. Each condition consisted of two practice sentences, followed by eight evaluation sentences. We only analyzed the evaluation sentences. For each condition, sentences were pulled randomly without replacement from the “mem1-5” set from the mobile Enron dataset (Vertanen and Kristensson, 2011). These sentences were in mixed case and included punctuation and sometimes numbers.
4.2.1. Independent variables and statistical tests.
We tested for statistical differences using a two-way repeated measures ANOVA. Our two independent variables were: 1) whether eye-tracking filtering was enabled (denoted Eye), and 2) whether participants typed with index fingers or all ten fingers (denoted Fingers). In the case of significant interaction effects, we tested for differences between conditions via Bonferroni corrected post-hoc tests. For Likert ratings, we used a non-parametric aligned rank transform (Wobbrock et al., 2011).
4.2.2. Participants and demographics.
We recruited 24 participants aged 18–22. Of these, 17 were male, 5 were female, and 2 preferred not to say. 23 participants strongly agreed with “I consider myself a fluent speaker of English”, and the remaining participant agreed with the statement. A total of seven participants reported AR experience, with four participants having used a HoloLens before, and three participants reporting use of phone AR apps (e.g. Pokémon Go).
@@ 4.3. Results
4.3.1. Entry rate.
We measured entry rate as in Experiment 1. Entry rates were 10.9 in Index, 9.4 in IndexEye, 7.1 in Ten, and 7.4 WPM in TenEye (Figure 8 left). There were significant main effects for both the use of eye-tracking and the number of fingers used (Table 2 left). As there was a significant interaction effect between the independent variables, we conducted post-hoc pairwise comparisons between conditions. All were significant except between Ten and TenEye.
@@
Entry rate (WPM) | Error rate (CER %) | Backspaces per final character | ||
Condition | Index | 10.9 1.4 [8.5, 13.9] | 0.63 0.82 [0.00, 3.11] | 0.12 0.06 [0.01, 0.27] |
IndexEye | 9.4 1.3 [6.9, 11.5] | 0.58 0.92 [0.00, 2.89] | 0.08 0.07 [0.02, 0.31] | |
Ten | 7.1 1.7 [3.1, 10.5] | 1.91 2.21 [0.00, 9.63] | 0.30 0.14 [0.07, 0.62] | |
TenEye | 7.4 1.9 [4.3, 11.0] | 1.27 1.54 [0.00, 6.52] | 0.16 0.11 [0.04, 0.39] | |
Independent | Eye | = , = , ¡ | = , = , = | = , = , ¡ |
variable | Fingers | = , = , ¡ | = , = , ¡ | = , = , ¡ |
Eye Fingers | = , = , ¡ | = , = , = | = , = , ¡ | |
Pairwise | Index IndexEye, ¡ | n/a | Index IndexEye, = | |
post-hoc | Index Ten, ¡ | Index Ten, ¡ | ||
tests | Index TenEye, ¡ | Index TenEye, = | ||
IndexEye Ten, ¡ | IndexEye Ten, ¡ | |||
IndexEye TenEye, ¡ | IndexEye TenEye, ¡ | |||
Ten TenEye, = | TenEye Ten, ¡ |
4.3.2. Error rate.
As in Experiment 1, we measured typing errors using character error rate (CER). The CER was 0.63% for Index, 0.58% for IndexEye, 1.91% for Ten, and 1.27% for TenEye (Figure 8 center). There was a significant main effect on the number of fingers used but not for the use of eye-tracking (Table 2 center). Thus it seems that regardless of the use of eye-tracking filtering, errors increased for ten finger typing.
@@
4.3.3. Correction rate.
To measure error corrections, we calculated the number of backspaces per final output character for each phrase. The number of backspaces pressed per output character was 0.12 for Index, 0.08 for IndexEye, 0.30 for Ten, and 0.16 for TenEye (Figure 8 right). There were significant main effects for both the use of eye-tracking and the number of fingers used (Table 2 right). As there was a significant interaction effect, we conducted post-hoc pairwise comparisons between conditions. All were significant with the exception of between Index and IndexEye, and between Index and TenEye. Notably, eye-tracking did significantly reduce the number of corrections required in ten-finger typing. This suggests that our filtering feature was successful at preventing some erroneous keystrokes.
4.3.4. Where participants looked.
@@
Figure 10 shows the average distance between the pressed key and eye gaze, before and after the key was pressed. In all conditions, the user looked most closely at the key at the instant it was pressed. Despite users tending to look more closely at the pressed key in the IndexEye and TenEye conditions, likely due to the restraints imposed by our eye-tracking feature, all four conditions followed the same pattern of eye movement. This supports our theory that users will naturally look where they intend to type. This indicates potential for incorporating this implicit gaze signal to improve a keyboard’s interface or algorithms (e.g. auto-correct or word predictions).
4.3.5. Finger usage.
We analyzed how often participants used each finger to type to measure how well participants followed our instructions to type using both hands. In the index finger conditions, participants generally used both index fingers and preferred using their right index finger. In the Index condition, on average, participants used their right index finger 62% of the time and ranged between 47% and 75% across participants. The results were similar in the IndexEye condition except that one participant used their right index finger exclusively even though we instructed all participants to try to use both hands. When we exclude that participant and average the others, we found that, similar to the Index condition, the right index finger was used 62% of the time and ranged from 53% to 78% across participants.
Boxplots showing which percentage taps were performed with each finger.
Figure 11 shows which fingers participants used in the ten-finger conditions. Participants used all ten fingers at least once in these conditions with only a few exceptions (four participants in Ten and one in TenEye). Typing with index fingers was preferred with the right index finger being used the most frequently. In the Ten condition, on average, participants typed with their left and right index fingers 17% and 30% of the time respectively. Interestingly, in the TenEye condition, this tendency was less pronounced at 12% and 21% for the left and right index fingers respectively. This shows that the eye-tracking feature may have increased participants’ confidence in typing with other fingers.
@@ 4.4. Subjective feedback
4.4.1. Likert statements.
After each condition, participants rated three statements on a 7-point Likert scale (Figure 12). Participants’ average ratings for “I entered text quickly” were: Index 4.4, IndexEye 3.9, Ten 3.1, and TenEye 3.3. The difference with respect to the number of fingers used was significant ( = ¡ ). The difference with respect to the use of eye-tracking was not significant ( = = ). There was no significant interaction effect ( = = ).
Participants’ average ratings for “I entered text accurately” were: Index 5.2, IndexEye 5.8, Ten 3.5, and TenEye 4.5. The differences were significant both with respect to the number of fingers () and the use of eye-tracking (). There was no significant interaction effect ().
Participants’ average ratings for “Entering text was physically easy and not tiring” were: Index 4.2, IndexEye 4.4, Ten 3.7, and TenEye 4.5. There was no significant difference for the number of fingers (), the use of eye-tracking (), or the interaction between them ().
4.4.2. Condition ranking.
We asked participants to rank the conditions and to explain their reasoning. The Index condition was most preferred with 16 of 24 participants ranking it first. Some representative comments were: “I didn’t have to worry about other fingers getting in the way”, “index fingers are both faster and more accurate for this”, and “I prefer the comfort of being able to relax my other fingers”. The Ten condition was never ranked as the first choice and the last choice by 15 participants.
@@
4.4.3. Likes and dislikes.
We coded participants’ likes and dislikes as in Experiment 1. Compared to the results of the condition ranking, these responses revealed a more nuanced result. 10 participants said they liked that Index only allowed two fingers but 7 disliked that limitation. After the Ten condition, 6 people said they liked being able to use all 10 fingers and 4 liked being able to use more fingers (but not all 10 fingers). Some representative comments were “[index fingers] were mentally easier because there was less error with less fingers able to hit random buttons” and ”the speed of typing with unrestricted 10 fingers was fun”. People generally disliked how hard it was to type in the Ten condition with 15 people reporting unexpected keypresses and 12 reporting poor tracking particularly for certain fingers. For example, one participant disliked “my fingers getting in the way while typing”.
Eye tracking appeared to make ten finger typing less error-prone. After TenEye, 11 participants liked how eye tracking helped and only one disliked the unexpected keypresses in the condition. Despite the great reduction in unexpected keypresses in TenEye, there were two dislikes of tracking in general and nine dislikes of tracking for certain fingers. For example, one participant said “middle and ring fingers don’t track properly”. Only one participant ranked TenEye as their top ranked choice but 17 ranked it above Ten. IndexEye was ranked as the first choice by seven participants (the second most popular choice after Index) and was ranked above IndexEye by 17 participants. After the IndexEye condition, four people disliked that they couldn’t use more fingers. The IndexEye and TenEye conditions each had four people who disliked being required to look at the spacebar. This suggests a possible interface improvement in which spacebar is excluded from the eye-tracking filtering.
4.4.4. AR improvements.
@@ 4.5. Discussion
Our goals in Experiment 2 were to 1) examine if users typed faster in midair using all ten fingers, and 2) to see if eye-tracking could reduce accidental key presses. We found users had a significantly higher entry rate in the Index and IndexEye conditions than in the Ten and TenEye conditions. Users also expressed an overwhelming preference for the index finger conditions. Users performed poorly in the Ten condition as evidenced by lower entry rates compared to the index finger conditions and more backspaces per output character then all other conditions. Participants frequently complained in open feedback about the hand-tracking. They also reported difficulty pressing keys using their ring fingers and frustration with a lack of responsiveness of the hand-tracking.
@@
Our hypothesis that users looked at the key they were attempting to press was supported by our eye-tracking data. We examined this by noting how often people were looking near the key they pressed in the Index and Ten conditions (conditions that did not use eye-tracking to disable keys). If we had disabled far away keys far as in IndexEye and TenEye, key presses would have been excluded 32.9% and 36.0% of the time in the Index and Ten conditions respectively. This suggests participants often looked where they were typing even when they were not required to do so. However, the IndexEye condition only excluded 10.4% of key presses. This notably lower amount may be simply because participants were more careful to match their gaze to the key they were pressing, lowering the number of unintentional key presses. However, in the TenEye condition, 36.7% of key presses occurred on disabled keys. This result is similar to the Ten condition and may show eye-tracking successfully filtered out unintentional key presses. However, it is important to note we do not know how many discarded key presses were due to unintentional presses being helpfully discarded versus erroneously discarded.
@@
A common complaint during the eye-tracking conditions was having to look at the spacebar. Since the spacebar was large and thumbs often do not move far from it when typing, participants felt this unnecessarily slowed typing. In the future, keeping the spacebar active regardless of gaze position may improve results. Another common complaint was that the keys were unresponsive or difficult to trigger. We are uncertain if this was due to hand-tracking issues or the particular threshold distance we used for triggering keys. Reducing this distance would have made it easier to trigger keys, potentially improving performance. However, a shorter distance could also increase accidental key presses. The half-second cooldown may have also limited text entry speed in some circumstance, but given all participants in our studies typed substantially slower than 24 WPM on average (i.e. two characters per second), we think this did not unduly slow typing. Additional work could investigate if different key actuation details could help improve performance.
There is also other evidence that eye-tracking was successful. Firstly, the addition of eye-tracking in TenEye significantly reduced the amount of backspaces per character (Figure 8 right). Secondly, participants rated the stated “I entered text accurately” significantly higher in the eye-tracking conditions. Overall, our results demonstrate that eye-tracking is an effective tool that should be considered when implementing ten-finger midair typing systems.
5. Limitations and Future Work
Our entry rates in both experiments were undoubtedly limited by the deterministic nature of our keyboard. It is likely users could have typed substantially faster if they only had to aim approximately within a key or so of their desired letters. Powerful auto-correct might even be able to sift through the many false key triggers typical of ten-finger midair typing. However, even in the case a MR keyboard supporting fast typing via auto-correct, it would still need a fallback mode to support users entering exactly what they want. This can be necessary for certain types of content, or to avoid an “auto-correct trap” in which a user’s desired and correct text is repeatedly auto “corrected” into incorrect text. Our work can help guide how an auto-correcting MR keyboard can best support fallback to deterministic input.
Past work on midair keyboards have commonly oriented keyboards either completely vertically (e.g. (Adhikary and Vertanen, 2021b; Markussen et al., 2013)) or tilted slightly away from vertical (e.g. (Adhikary and Vertanen, 2021a; Dudley et al., 2018)). However, in most past studies, participants were standing and not seated as in our study. We opted instead for a 15 degree tilt from horizontal for both our studies. This felt more comfortable in pilot testing and more closely matches users’ real-world typing experience on desktop keyboards which often have an upward tilt. A similar tilt of 20 degrees (Dudley, 2025) was used in another study which explored seated typing on or near a surface (Dudley et al., 2019). While this tilt appeared to work well for our participants, further exploration of user performance and preference with regards to keyboard tilt could be useful.
@@ 5. Limitations and Future Work
Our participant population was young, tech savvy, and mostly male. While we suspect the differences we found between conditions were driven by challenges related to midair target selection and hand-tracking, we caution results could vary for other populations.
An interesting avenue for future research would be to use auxiliary sensors (e.g. the AR headset’s microphone or a smartwatch) to detect taps on a physical surface instead of, or in addition to, finger collision with the keys. This could improve tap detection accuracy of virtual targets positioned on physical surfaces. In particular, the tactile feedback of actual contact with a surface could mitigate the problem of unintentional key presses in ten-finger typing.
While we a current commodity headset struggled to track all ten fingers, we found most participant complaints mostly revolved around certain fingers such as the ring and middle finger. An interesting future study might explore typing with a more limited set of fingers (e.g. the index fingers plus thumbs and pinky fingers). This could allow users to use some fingers for common keys (e.g. thumbs for spacebar and pinky finger for backspace).
In Experiment 2, we used gaze location to make a hard decision about which keys were active. This threshold was determined based on pilot testing and given the eye-tracking accuracy of our headset. It is possible a different threshold, or a more accurate eye-tracker, would improve performance of this feature. It would also be interesting to investigate improving location or tap detection accuracy via a softer, probabilistic fusion of the hand and eye-tracking signals. We have included in our supplementary materials666https://osf.io/q4n72 detailed participant log files from both experiments. This includes when participants triggered individual keys as well as continuously sampled observations from the hand and eye tracker for Experiment 2. This data could allow exploring potential gains from incorporating such changes into future midair keyboards. We have also included the source code for the keyboards we used in both experiments.